OpenStudy (anonymous):

A ball accelerates uniformly at +4.0m/s^2. If the ball was initially moving at +0.20m/s and the average velocity was +4.2 m/s, then how long will it take to reach the final velocity? Please show how you get this. I am sooo lost. I want to learn how, not just get the answer. Thank you!

OpenStudy (anonymous):

You are given initial velocity and the average velocity. Average velocity = (final velocity - initial velocity)/2 so you can solve for final velocity. 4.2 = (x - 0.2)/2 8.4 = x-0.2 x = 8.6 m/s so final velocity is 8.6 m/s. Final velocity = initial velocity + acceleration(time) You might see this as Vf = Vi + at 8.6 = 0.2 + 4.0(t) Time = 2.1 seconds